geerky42 (geerky42):

\[\int_0^{100} \sqrt{1+(f'(x))^2}dx\]Where \(f(x) = 75\left( e^{\dfrac{x}{150}} + e^{-\dfrac{x}{150}} \right)\)

OpenStudy (anonymous):

\[\int_0^{100} \sqrt{1+(f'(x))^2}dx\]Where \(f(x) = 75\left( e^{\dfrac{x}{150}} + e^{-\dfrac{x}{150}} \right)\) Just copying the original question, because OS doesn't show \(\LaTeX\) for me in the opening post. So it's easier to read.

OpenStudy (anonymous):

First calculate f'(x) . Can you do that?

hartnn (hartnn):

f(x) =.... ? hint : hyperbolic functions.

geerky42 (geerky42):

I end up with \(\displaystyle\int_0^{100}\sqrt{e^{x/75}+ e^{-x/75}-2}\space dx\) I don't know what to do next... Any idea? I don't know hyperbolic functions well.

hartnn (hartnn):

look for the definition of cosh x

hartnn (hartnn):

this will greatly simplify your integral

geerky42 (geerky42):

Got to go... We can discuss later. Sorry!

OpenStudy (anonymous):

It would go like this : \(f'(x) = 75(e^{\cfrac{x}{150} }\times \cfrac{1}{150} - \cfrac{1}{150} \times e^{-\cfrac{x}{150}})\) \(f'(x) = \cfrac{-1}{2} (e^{\cfrac{x}{150}} - e^{-\cfrac{x}{150}})\) why don't put f'(x) = tan t

OpenStudy (anonymous):

Go to go now. Sorry would look later.